OpenStudy (anonymous):

Find slope intercept equation for the line with the indicated slope Slope 5/4 and the y intercept (0,-8)

sam (.sam.):

\[\huge y=mx+c\] \[\huge -8=\frac{5}{4}(0)+c\] \[\huge c=-8\] \[\huge y=\frac{5}{4}x-8\]